# A Thorough Assessment of the Non-IID Data Impact in Federated Learning

### v1

@@ IV Experimentation Setup
This work considers six widely employed datasets to train the CL and FL models. Four of them, i.e., CIFAR10 [19], FMNIST [46], Physionet 2020 [14], Covtype [4], serve to simulate label, feature, and quantity skew, while the remaining two datasets, i.e., 5G Network Traffic flows [8], Snapshot Serengeti [41] work for spatiotemporal skew simulation. Table I summarizes the main characteristics of the datasets.
Table I: Characteristics of the datasets
@@
Dataset Type #training examples #test examples #features # classes Class distr.
@@
Models
We adopt a well-studied convolutional neural network (CNN) broadly applied in computer vision [5] for the image datasets. It includes one input layer and three convolutional blocks—where the first two blocks each have a convolutional layer followed by a max pooling layer, and the final block contains a convolutional layer and a flattened layer. The initial convolutional layer has 32 filters, whereas the subsequent two layers each have 64 filters, all with a $3\times 3$ filter size and ReLu as the activation function. In the dense section of the network, there is one dense layer with 64 neurons using ReLu as the activation function. Additionally, we employ in our tests the transfer learning models EfficientNetB0 [42] and MobileNetV2 [39] since they produce higher classification power results for the datasets studied.
For the tabular datasets, we use a deep neural network (DNN), selected because it is widely employed in classification tasks with tabular data [37]. It comprises one input layer, three hidden layers, and one output layer [32]. The input layer uses as many units as the number of features used in the training set. The three layers contain 500 hidden units each, and the last layer is formed by considering the neurons equal to the number of classes to predict. Additionally, the hidden layers used the ReLu activation function and the output layer a SoftMax function. We use Adam as our optimizer with a learning rate of 0.001 for $K=30$ clients and a batch size of 64. We train the models for 40 communication rounds and ten local epochs for all aggregation algorithms except MOON, which is trained for 100 communication rounds. We run the experiments on all six datasets using five or ten different data partitions (i.e., five or ten random seeds).
IV-1 Hyperparameters tuning
@@
IV-A Methods for Synthetic Partitioning
The datasets introduced previously are all centralized, and we must split them while controlling the level of non-IIDness to train FL models. For such a purpose, we use the FedArtML [17] tool to partition each dataset among the clients. The following lines explain the partition method employed to simulate each skewness type.
IV-A1 Label Skew
We use the Dirichlet distribution (DD) to partition data among clients based on label distribution. The DD generates random numbers summing to one, controlled by parameter $\alpha$. Higher $\alpha$ values (e.g., 1000) create similar local distributions, while lower values increase the chance of clients having examples from a single, randomly chosen class [26]. Notice that the DD is the multivariate generalization of the Beta distribution, and the latter is the generalization of the Uniform distribution. Therefore, the partition of the datasets using DD is a skewed split of the data distribution [25].
Figure 2: Distribution of CIFAR10 among 30 clients for different levels of non-IIDness.
We quantify the level of non-IIDness of the data partitions on the clients using the HD. Our levels of non-IIDness, quantified using HD, would be approximately equal to the following values: $\{0.0,0.25,0.5,0.75,0.9\}$. Fig. 2 exemplifies the partition distribution for label skew using thirty clients. All ten classes get evenly distributed among every client in the IID scenario ($\alpha=1000,\textsf{HD}=0.0$). As we increase the $\alpha$ parameter in the DD, the distribution of classes among clients becomes more diverse. In the extreme case of $\alpha=0.03,\textsf{HD}=0.9$, certain classes are absent in some local nodes.
IV-A2 Feature Skew
For simulating feature skew, we employed two diverse methods from FedArtML to test their properties:
Hist-Dirichlet-based method: The process starts by characterizing the attributes of each client using other average values and then subjecting them to a binning procedure. Subsequently, it establishes the participation of each feature category within each client using the DD with a specified $\alpha$. Unlike the Gaussian Noise approach, this method distributes the data among the clients without modifying the features.
We measure the non-IIDness with the HD among the features across clients (FHD) within the range $\{0,0.25,0.5,0.75,0.9\}$.
IV-A3 Quantity Skew
We employ the MinSize-Dirichlet method, which specifies the DD’s $\alpha$ value and generates the desired participation proportions for each client. Subsequently, a minimum required size, which we refer to as “the minimum number of examples,” is established for each client. Thus, the minimum proportion size, denoted as $MinSize$, is calculated as $\textit{MinSize}=\frac{\textit{MinRequiredSize}}{n}$, where $n$ represents the total number of examples in the centralized dataset. If the designated proportions fall below MinSize, it substitutes them with MinSize. Finally, the proportions are normalized to fall from 0 to 1.
We assess the level of non-IIDness using the HD for quantity skew (QHD) within the range $\{0,0.10,0.17\}$.
IV-A4 Spatiotemporal (SPT) skew
The primary constraint in this partition process is that the dataset must contain a categorical variable of space (i.e., places, cities, latitude, longitude, etc.) or time (i.e., time, months, years, etc.) to use as the partitioning variable. For instance, Fig. 3 depicts the distribution of labels along the date of the 5GNTF dataset. In this case, the space variable employed to create the federated data is the flow’s date expressed in year-month-day format (categorical).
We use the St-Dirichlet method, which employs the DD to segment the data based on spatial (SP skew) or temporal (TMP skew) categories to distribute the data among federated clients. We assess the level of non-IIDness using the HD for spatiotemporal skew (STHD) within the range {0, 0.25, 0.5, 0.75, 0.9}.
IV-A5 Hardware Specification
@@
We used an Ubuntu 22.04.4 LTS machine with 200 GB of disk, Intel(R) Xeon(R) Platinum 8259CL CPU @ 2.50GHz processor, 16 processors, 125 GB of RAM, and Python 3.10.12 to run the experiments. The FL models were trained using the Flower [2].
V Results and Highlights
This section depicts the main highlights and results of the experiments defined in Section IV. Each result gets grouped by label, feature, quantity, and spatiotemporal.
V-A Label Skew
@@
This subsection examines how label skewness in the client data affects the models’ performance. Consider that the Covtype is an unbalanced dataset regarding the labels, and CIFAR10, FMNIST, and Physionet are balanced datasets. The accuracy of the models on each centralized dataset is our baseline for comparing the accuracy of the models generated in FL.
V-A1 Classification power
@@
In this subsection, we focus on the findings from the simulations to compare different aggregation algorithms and datasets regarding their classification power (a.k.a accuracy).
@@
Table II: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IIDness as measured by HD for $K=30$. Each model has undergone ten different trials.
@@
Number of times that performed the best 3 1 8 2 6
@@
Highlight 1: The drop in the model’s performance appears in a double threshold. When HD is bigger than 0.5 and 0.75, a notable performance decline becomes immediately evident.
@@
Previous works claim that the data non-IIDness affects the performance of FL models [27, 29, 16]. Nevertheless, for the first time, we demonstrate that the effect of non-IIDness is not the same under all the levels of heterogeneity. Fig. 4 depicts the accuracy change by varying the level of non-IIDness of the data distributions among the clients measured by HD concerning the baseline model created in the centralized setting. As the non-IIDness of the data partitions increases, the model’s accuracy decreases. When the HD between data distributions of the clients exceeds 0.75, the drop becomes more drastic compared to previous levels.
@@
Figure 4: Changes in the models’ accuracy considering different levels of non-IIDness measured by HD for $K=30$.
Highlight 2: For high levels of non-IIDness the performance of transfer learning models gets more degraded
@@
Figure 5: Changes in the models’ accuracy considering different levels of non-IIDness measured by HD for $K=30$ using CNN, EfficientNetB0 and MobilbeNetV2 on CIFAR10 dataset.
Fig. 5 illustrates the accuracy of models created using three different architectures: the CNN discussed earlier, EfficientNetB0, and MobileNetV2, both of which utilize transfer learning. The figure clearly shows a decline in model performance at the double thresholds after 0.5 and 0.75 HD. Notably, in pathological non-IID scenarios, the transfer learning models perform worse than our naive CNN model. This behavior can be attributed to the fact that only the last layers’ weights are updated locally, leading to skewed model weights and reducing the effectiveness of the aggregated model in capturing the overall dataset’s features, making transfer learning less suitable in such cases.
@@
Highlight 3: Aggregation algorithms such as Rand, POC, and MOON are particularly vulnerable to performance degradation under conditions of high non-IIDness.
@@
Consider the case when the HD is 0.9 (i.e., high non-IIDness) for all the datasets presented in Table II. When comparing the accuracy of Rand and POC versus their correspondent CL performance, those algorithms tend to have a higher drop in performance than the other aggregation methods. In FL scenarios with high non-IIDness, the distribution of data labels across clients is highly uneven. The latter means that specific clients may have more or different data types than others. Under such a scenario, Rand and POC may inadvertently select clients with skewed or unrepresentative data distributions, leading to poor generalization performance when aggregating their updates. MOON has the sharpest decrease in performance when the paradigm switches from CL to FL. Penalizing the current model with past and global representations made by each client will not help the model attain improved performance.
@@
Table III: The performance’s decrease range of the model for the four datasets, moving from the lowest (HD=0) to the highest (HD=0.9) levels of non-IIDness and considering FedAvg.
@@
Consider the IID case (i.e., HD = 0) and the most extreme non-IID case (i.e., HD = 0.9) for each dataset as depicted in Table III. Notice that the range of decrease in the reported performance (accuracy of HD=0 - accuracy of HD=0.9) is higher for the unbalanced dataset (Covtype) than for balanced datasets (CIFAR10, FMNIST, and Physionet). The more significant decrease in performance for unbalanced datasets can relate to the challenge of learning from skewed and less representative data. In non-IID scenarios, where clients have different class distributions, models trained on these unbalanced datasets may struggle to generalize well across clients.
V-A2 Convergence

### v2

@@ 4 Experimentation Setup
This work considers eight widely employed real datasets to train the centralized learning (CL) and FL models. Four of these, i.e., CIFAR10 [36], FMNIST [72], Physionet 2020 [26], and Covtype [8] serve to simulate label, feature, and quantity skew. To further examine label skew in scenarios with a considerably larger number of classes, we also included CIFAR100 [35]. The remaining three datasets, i.e., 5G Network Traffic flows [14], MHEALTH [5], and Snapshot Serengeti [62], are used to simulate spatiotemporal skew. Table 2 provides an overview of the main characteristics of each dataset.
Table 2: Characteristics of the datasets
@@
Dataset Type #training examples #test examples #features #classes Classes distribution
@@
MHEALTH Tabular 851,021 364,724 14 13 Unbalanced
Models.
We adopt a well-studied CNN broadly applied in computer vision [10] for the CIFAR10 and FMNIST datasets. It includes one input layer and three convolutional blocks, where the first two blocks each have a convolutional layer followed by a max pooling layer, and the final block contains a convolutional layer and a flattened layer. The initial convolutional layer has 32 filters, whereas the subsequent two layers each have 64 filters with a $3\times 3$ filter size and ReLu as the activation function. In the dense section of the network, there is one dense layer with 64 neurons using ReLU as the activation function. We utilized a ResNet9 for the CIFAR100 dataset. Additionally, we employed in our tests the transfer learning models EfficientNetB0 and MobileNetV2 since they produce higher classification power results for the datasets studied.
For the tabular datasets, we use a DNN, selected because it is widely employed in classification tasks with tabular data [58]. It comprises one input layer, three hidden layers, and one output layer [50]. The input layer uses as many units as the number of features in the training set. The three layers contain 500 hidden units each, and the last layer is formed by considering the neurons equal to the number of classes to predict. Additionally, the hidden layers used the ReLu activation function, and the output layer used a SoftMax function. We use Adam as our optimizer with a learning rate of 0.001 for $K=30$ clients and a batch size of 64. In our simulations, models were trained for 40 communication rounds and 10 local epochs, except for those using the MOON aggregation algorithm and those involving the CIFAR100 dataset, which were trained for 100 communication rounds to ensure convergence in those settings. To ensure reproducibility and statistical validity, we executed all experiments across the datasets using five and ten distinct data partitions generated from fixed random seeds.
4.1 Hyperparameters Tuning
@@
  • FedAvg: We do not set any specific tuning process for this algorithm [47].
@@
4.2 Hardware Specification
@@
We used an Ubuntu 22.04.4 LTS machine with 200 GB of disk, Intel(R) Xeon(R) Platinum 8259CL CPU @ 2.50GHz processor, 16 processors, 125 GB of RAM, and Python 3.10.12 to run the experiments. The FL models were trained using the Flower [6] platform.
4.3 Performance Metrics
This subsection describes the performance and convergence metrics considered in our experiments and a justification for their use.
Accuracy [64, 69, 49]. It refers to the proportion of correctly classified instances compared to the total data size. Higher values of accuracy indicate a better model performance. It can be calculated as follows:
$Acc=\frac{\sum_{k=1}^{K}C_{k}}{\sum_{k=1}^{K}n_{k}}$ (3)
where $C_{k}$ indicates the number of correctly classified samples on client $k$ and $n_{k}$ is the number of data samples on client $k$. We performed five and ten independent trials using different random seeds for the experiments. To ensure robust and reliable results, we report the mean accuracy and the standard deviation across these trials, providing a comprehensive view of the model’s performance variability.
Curvature [17, 24]. We incorporate curvature as a metric to identify points along the accuracy curve with respect to HD where model performance changes considerably. Given a parametric curve $\alpha(t)=(x(t),y(t))$, where $x(t)$ denotes the level of non-IID data quantified by HD and $y(t)$ the corresponding model accuracy, the curvature $\kappa(t)$ at that point is defined as:
$\kappa(t)=\frac{x^{\prime}(t)y^{\prime\prime}(t)-x^{\prime\prime}(t)y^{\prime}(t)}{\left(x^{\prime}(t)^{2}+y^{\prime}(t)^{2}\right)^{3/2}}$ (4)
where $x^{\prime}(\cdot)$ denotes the first-order derivative and $x^{\prime\prime}(\cdot)$ is the second-order derivative.
Number of times detected as critical point (#Detected as critical point). To identify critical points related to the effects of non-IID data, we count how many times each HD value is detected as critical based on curvature. Specifically, points where the curvature satisfies $\kappa\geq 1$ are considered indicators of sharp performance degradation. This count-based metric highlights HD values where degradation occurs consistently across the different models built by varying the label skew of the clients’ data distributions, reflecting the robustness and consistency of a critical point of change across different settings.
Average curvature. This metric captures the overall sharpness of performance change at each HD value by averaging curvature values across different models built under the same non-IID data. By averaging out the curvature values across models, it offers a more robust and reliable estimate of how critical each point truly is. A higher average curvature indicates a sharper and more consistent performance drop, pointing to greater sensitivity or instability at that level of non-IID data. This metric aims to identify the points where performance degradation not only present but also most pronounced, helping us more effectively evaluate and compare the flagged HD values (0.5 and 0.75) as critical points.
Number of times that performed the best [38]. This metric quantifies how frequently an aggregation algorithm performs better than other approaches across multiple experimental trials. A higher count indicates greater consistency and robustness. This metric helps identify which methods consistently excel across different data partitions, which is particularly valuable in FL, where non-IID data distributions can lead to high-performance variance between trials.
Rounds-to-accuracy (RTA) [71]. This metric measures the minimal number of communication rounds needed for the global model to achieve at least 90% of the maximum accuracy among the aggregation algorithms. It reflects the efficiency of the FL process since lower values indicate a faster model’s convergence.
5 Label Skew Results
@@
This section examines how label skew in the client data affects the models’ performance. Consider that the Covtype is an unbalanced dataset regarding the labels, and CIFAR10, FMNIST, CIFAR100, and Physionet are balanced datasets. The accuracy of the models on each CL is our baseline for comparing the accuracy of the models generated in FL.
5.1 Synthetic Partitioning Method
Using the FedArtML tool [33], we employed the Dirichlet distribution (DD) to partition data among clients based on label distribution. The DD generates random numbers summing to one, controlled by the parameter $\alpha$. Higher $\alpha$ values (e.g., 1000) create similar local distributions, while lower values increase the chance of clients having examples from a single, randomly chosen class [43]. The selected values {1000, 6, 1, 0.3, 0.03} allow us to examine the impact of varying degrees of non-IID data on FL performance. Notice that the DD is the multivariate generalization of the Beta distribution, and the Beta distribution is itself a generalization of the Uniform distribution.. Therefore, the partition of the datasets using DD is a skewed split of the data distribution [42].
Figure 2: Distribution of CIFAR10 among 30 clients for different levels of non-IID data. The x-axis shows distinct $\alpha$ values used to partition the data and the resulting HD for clients from 1 to 30. The y-axis shows the participation of each class depicted on the z-axis.
We quantify the degree of non-IID data across clients using the HD. Specifically, we partition the data using the DD’s $\alpha$ values $\{1000,6,1,0.3,0.03\}$ to achieve distinct HD levels $\{0.0,0.25,0.5,0.75,0.9\}$ to cover a representative spectrum of non-IID data ranging from fully IID to highly non-IID partitions. For instance:
  • A DD concentration parameter $\alpha=1000$ yields IID data ($\textsf{HD}\approx 0.0$), as labels are uniformly distributed.
  • Conversely, $\alpha=0.03$ produces highly skewed partitions ($\textsf{HD}\approx 0.9$).
Thus, each $\alpha$ value maps to a unique HD based on the label distribution, enabling controlled experimentation across non-IID scenarios. Figure 2 exemplifies the partition distribution for label skew using thirty clients. All ten classes get evenly distributed among every client in the IID scenario ($\alpha=1000,\textsf{HD}=0.0$). As we increase the $\alpha$ parameter in the DD, the distribution of classes among clients becomes more diverse. In the extreme case of $\alpha=0.03,\textsf{HD}=0.9$, certain classes are absent in some clients.
5.2 Classification Power
@@
In this subsection, we focus on the findings from the simulations to compare different aggregation algorithms and datasets regarding their classification power (a.k.a. accuracy).
@@
Table 3: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IID data as measured by HD for $K=30$. Each model has undergone ten different trials (random seeds).
@@
Number of times that performed the best 4 1 12 2 6
Table 4: Curvature values derived from the CNN model’s accuracy trends at various levels of non-IID data, highlighting points of sharp change in performance under increasing label skew across five datasets.
Dataset HD=0.25 HD=0.50 HD=0.60 HD=0.65 HD=0.70 HD=0.75 HD=0.80 HD=0.85
CIFAR10 0.2 0.3 1.1 0.5 1.6 3.9 4.4 2.8
Covtype 0.3 0.4 0.2 0.3 0.2 3.5 6.6 2.9
FMNIST 0.0 0.0 0.5 0.6 0.9 1.3 1 3.3
Physionet 0.1 0.2 0.5 2.2 1.3 1 3.3 3.2
CIFAR100 0.0 0.2 0.1 0.1 1.2 1.5 2.5 1.8
#Detected as critical point 0 0 1 1 3 5 5 5
Average curvature 1.2 0.2 0.5 0.75 1 3.7 3.57 2.8
Table 5: Curvature values derived from the accuracy trends of CNN, EfficientNetB0, and MobileNetV2 models on the CIFAR-10 dataset at various levels of non-IID data, indicating points of sharp performance change.
Model HD=0.25 HD=0.50 HD=0.60 HD=0.65 HD=0.70 HD=0.75 HD=0.80 HD=0.85
CNN 0.2 0.3 1.1 0.5 1.6 3.9 4.4 2.8
EfficientNetB0 0.2 0.5 2.3 3.1 4.7 3.7 2.1 0.7
MobileNetV2 0.2 0.3 2.0 2.9 5.5 4.8 1.1 0.3
#Detected as critical point 0 0 3 2 3 3 3 1
Average curvature 0.2 0.4 1.8 2.2 3.9 4.1 2.5 1.3
@@
Highlight 1: The drop in the model’s performance for label skew appears in a double threshold. A notable performance decline is immediately evident when the HD exceeds 0.5 and 0.75.
@@
Previous works claim that non-IID data affects the performance of FL models [44, 45, 31]. Nevertheless, for the first time, we showcase that the effect of non-IID data is not the same under all levels of heterogeneity. Figure 3 depicts the accuracy change by varying the level of non-IID data distributions among the clients measured by HD concerning the baseline model created in the centralized setting. As the non-IID data partitions increase, the model’s accuracy decreases. When the HD between data distributions of the clients exceeds 0.75, the drop becomes more drastic compared to previous levels.
@@
Figure 3: Changes in the models’ accuracy considering different levels of non-IID data measured by HD for $K=30$.
To more precisely characterize the inflection points in model performance, we compute the curvature of the accuracy curves across multiple datasets. As shown in Table 4, curvature values highlight sharper changes at HD = 0.75 and beyond. Notably, models also start to experience a sharper decline in accuracy after HD = 0.5, indicating the beginning of instability. The aggregated metric "#Detected as critical point" shows that HD = 0.75 is identified as a critical point in all five datasets, and the average curvature at this point (3.7) is the highest across all HD values.
One possible explanation for this double-threshold effect is that when HD surpasses 0.5, the model starts experiencing a noticeable decline due to increasing divergence in local distributions, leading to a degradation in the global model’s generalization. However, beyond HD = 0.75, the level of heterogeneity may reach a critical point where client models become overly specialized to their local data, significantly reducing the effectiveness of global aggregation. This sharp accuracy drop suggests that at extreme levels of non-IID data, FedAvg struggles to find a well-generalized solution, potentially due to conflicting optimization directions from highly dissimilar client updates.
Highlight 2: Transfer learning models exhibit greater sensitivity to variations in clients’ label distributions, with performance degrading more rapidly and sharply as non-IID data increases.
Figure 4 illustrates the accuracy of models created using three different architectures: the CNN discussed earlier, EfficientNetB0, and MobileNetV2, both of which utilize transfer learning. As HD increases, all models experience performance degradation. Transfer learning models exhibit a more pronounced decline under high non-IID data compared to the CNN which stems from their reliance on frozen feature extractors, limiting adaptation to heterogeneous data. As HD increases, local updates to the final layers create misaligned feature representations, reducing the effectiveness of global aggregation. In contrast, the CNN trained from scratch better adapts to decentralized data, making it more robust in extreme non-IID settings.
@@
Figure 4: Changes in the models’ accuracy considering different levels of non-IID data measured by HD for $K=30$ using CNN, EfficientNetB0, and MobileNetV2 on the CIFAR10 dataset.
Table 5 presents the curvature analysis for the same models. The curvature values for the transfer learning models not only confirm the sharper performance drops but also reveal that these declines begin earlier along the HD spectrum. The peaks in curvature, which quantify the steepness of accuracy degradation, occur sooner and with higher intensity compared to those in Table 4. This suggests that transfer learning models are more sensitive to label distributional shifts over the clients’ data. Still, the presence of distinct inflection points after HD = 0.5 and at more prominent at HD = 0.75 further supports the existence of a double threshold pattern in performance degradation under non-IID conditions.
@@
Highlight 3: Aggregation algorithms such as Rand, POC, and MOON are particularly vulnerable to performance degradation under conditions of high non-IID data.
@@
Consider the case when the HD is 0.9 (i.e., high non-IID data) for all the datasets presented in Table 3. When comparing the accuracy of Rand and POC versus their corresponding CL performance, those algorithms tend to have a higher drop in performance than the other aggregation algorithms. In FL scenarios with high non-IID data, the distribution of data labels across clients is highly uneven. It means that specific clients may have more or different data types than others. Under such a scenario, Rand and POC may inadvertently select clients with skewed or unrepresentative data distributions, leading to poor generalization performance when aggregating their updates.
MOON has the sharpest decrease in performance when the paradigm switches from CL to FL. MOON’s contrastive loss, designed to align local and global representations, becomes ineffective when client distributions are too divergent, as past representations no longer serve as meaningful anchors. This misalignment exacerbates performance degradation, making these methods less suited for extreme non-IID scenarios.
@@
Consider the IID case (i.e., HD = 0) and the most extreme non-IID case (i.e., HD = 0.9) for each dataset as depicted in Table 6. Notice that the range of decrease in the reported performance (accuracy of HD=0 - accuracy of HD=0.9) is higher for the unbalanced dataset (Covtype) than for balanced datasets (CIFAR10, FMNIST, CIFAR100, and Physionet).
Table 6: The performance’s decrease range of the model for the four datasets, moving from the lowest (HD=0) to the highest (HD=0.9) levels of non-IID data and considering FedAvg.
@@
The sharper performance drop in unbalanced datasets under high non-IID data derives from the compounding effects of class imbalance and non-IID data. In such cases, certain classes may be overrepresented in specific clients while being nearly absent in others, leading to biased local models. These biased updates fail to capture the overall class distribution when aggregated, resulting in poor generalization. In contrast, balanced datasets distribute class information more evenly across clients, mitigating this effect and leading to a less severe performance decline.
5.3 Convergence